Q7) Find the general solution for each of the following equations a) y +44t5y=o b) учу – 2 = + Чg + цо y
Answer #1

a y +44 + 5y zo Let D = d dre (D+42+5) Y=0 = 7 D = - 4 + √16-20 Q -uta -gt then general S8 69 is C.F +P.I - 2x G(50) = (c,@ o ر ا ا कर ) tyD 1.6 yll tuylthy + blyso Auxildan eh Pult Dom m²+um+y =) m = -2;-2 Soph are general -g2 yang = (c, +ga) e